Having problems with this question:
Find the coordinates of the vertex for the parabola given by the equation f(x)=-3x^2-x-1.

Answer by Alan3354(69443)

Find the coordinates of the vertex for the parabola given by the equation
f(x)=-3x^2-x-1.
----------
The vertex is on the LOS, Line of Symmetry, or the Axis of Symmetry
-----
The LOS is x = -b/2a
x = 1/-6 = -1/6
----
f(-1/6) = -1/12 + 1/6 - 1 = -11/12
--> vertex (-1/6,-11/12)